OpenStudy (anonymous):

When adding 2pi to get a positive coterminal angle, is it only added to the numerator... eg. pi/6 + 2pi = 2pi/6?

OpenStudy (anonymous):

I believe it would be pi/6 + 2pi/1 so lcd would be 6. which then becomes pi/6 +12pi/6= 13pi/6

OpenStudy (aum):

\[ \frac{\pi}{6} + 2\pi = \pi\left(\frac 16 + 2\right)= \pi\left(\frac 16 + \frac{12}{6} \right) = \pi\frac{13}{6} = \frac{13}{6} \pi \]